OpenStudy (anonymous):

If the base of a square pyramid is 6 centimeters on each side, and it has a volume of 396 cubic centimeters, what is the height of the pyramid

OpenStudy (anonymous):

volume = 1/3 side^2 * height 396 = 1/3 6^2 * height